# Three-dimensional Imaging of Pion using Lattice QCD: Generalized Parton Distributions

### v1

@@ B.1 Matching coefficient
The NNLO matching kernel with LRR and RGR can be expanded to ${\cal O}(\alpha_{s})$ as [58, 57]
@@
$\quad+\frac{\mu_{0}}{yP_{z}}\delta{\cal C}_{0,{\rm LRR\;rest}}{\cal C}_{\rm hybrid\;LRR}+{\cal O}(\alpha_{s}^{3}),$
@@
where $C_{F}=4/3$, $C_{A}=3$, $T_{F}=1/2$, and $n_{f}=3$ for our lattice ensemble. The 1-loop and 2-loop matching coefficients ${\cal C}^{(1)}$ and ${\cal C}^{(2)}$ are constructed as
${\cal C}^{(1)}={\cal C}^{(1)}_{\rm LL}+{\cal C}^{(1)}_{\rm hybrid\;LL}+{\cal C}^{(1)}_{{\rm NLL}}\ln\left(\left(\frac{\mu_{0}}{2P_{z}}\right)^{2}\right),$ (29)
${\cal C}^{(2)}={\cal C}^{(2)}_{\rm LL}+{\cal C}^{(2)}_{\rm hybrid\;LL}+\left({\cal C}^{(2)}_{{\rm NLL}}+{\cal C}^{(2)}_{{\rm hybrid\;NLL}}\right)\ln\left(\left(\frac{\mu_{0}}{2P_{z}}\right)^{2}\right)+{\cal C}^{(2)}_{{\rm NNLL}}\left[\ln\left(\left(\frac{\mu_{0}}{2P_{z}}\right)^{2}\right)\right]^{2},$ (30)
where ${\cal C}_{\rm LL/NLL/NNLL}$ represent the normal LL/NLL/NNLL terms, ${\cal C}_{\rm hybrid\;LL/NLL}$ denote the LL/NLL terms in the hybrid scheme. In the last line of Eq. (23), ${\cal C}_{\rm hybrid\;LRR}$ is the LRR term in the hybrid scheme [58, 57]
${\cal C}_{\rm hybrid\;LRR}$ $=0.575\mu_{0}\Bigg{\{}\frac{e^{-A}p_{z}(1+A+A^{2})}{\epsilon_{m}^{2}\pi}+\frac{1}{\pi}\Bigg{[}\frac{e^{-A}z_{s}\sin\left[Bz_{s}p_{z}\right]}{B}+\frac{e^{-A}p_{z}}{\left[\epsilon_{m}^{2}+p_{z}^{2}B^{2}\right]^{2}}$ (31)
×$\times\Big{(}\cos\left[Bz_{s}p_{z}\right]\left(\epsilon_{m}^{2}-B^{2}p_{z}^{2}+\epsilon_{m}^{3}z_{s}+\epsilon_{m}p_{z}^{2}B^{2}z_{s}\right)-\sin\left[Bz_{s}p_{z}\right]\left(2\epsilon+B^{2}p_{z}^{2}z_{s}+\epsilon_{m}^{2}z_{s}\right)Bp_{z}\Big{)}\Bigg{]}\Bigg{\}}_{+},$
where $e^{-A}=e^{-\epsilon_{m}z_{s}}$ is a small exponential decaying factor, $B=1-\frac{x}{y}$, $p_{z}=yP_{z}$, and {$\{\cdots\}_{+}$ means a plus function. Additionally, $\delta{\cal C}_{0,{\rm LRR\;rest}}$ represents the rest LRR term
$\delta{\cal C}_{0,{\rm LRR\;rest}}=-0.575\kappa$$\Bigg{\{}\alpha_{s}+\frac{\alpha_{s}^{2}\left[2\beta_{0}^{2}+\beta_{1}(1+c_{1})\right]}{4\pi\beta_{0}}+e^{-\frac{2\pi}{\alpha_{s}\beta_{0}}}\left[4\pi\beta_{0}+\alpha_{s}c_{1}\beta_{1}+\frac{\alpha_{s}^{2}\beta_{1}c_{2}(\beta_{1}-2\beta_{0}^{2})}{4\pi\beta_{0}}\right]$ (32)
×$\times\frac{1}{2\beta_{0}^{2}}\Re\left[E_{1+\frac{\beta_{1}}{2\beta_{0}^{2}}}\left(-\frac{2\pi}{\alpha_{s}\beta_{0}}\right)\right]\Bigg{\}},$
where $E_{n}(z)=\int_{1}^{\infty}e^{-zt}/t^{n}{\rm d}t$ is the generalized exponential integral, $\Re$ denotes the real part, and
@@
with $b=\beta_{1}/(2\beta_{0}^{2})$.
B.2 Evolution coefficient
@@
where $a_{s}=\alpha_{s}(\mu_{0})/(4\pi)$, $t=\ln(\mu^{2}_{0}/\mu^{2})$, ${\cal P}^{(0)}_{qq}$ is the 1-loop splitting function, and ${\cal P}^{V(1)}_{qq}$ and ${\cal P}^{V(2)}_{qq}$ are the 2-loop and 3-loop splitting functions for the valence quark [81, 82]. They are given by
@@
where the definitions of $P_{F}(z)$, $P_{G}(z)$, $P_{n_{f}}(z)$ and $P_{A}(z)$ can be found in Ref. [81]. As for ${\cal P}^{V(2)}_{qq}$, we take the approximate solution ${\cal P}^{V(2)}_{qq}(z)=P_{\rm ns}^{(2)-}(z)$ in Eq. (4.23) of Ref. [82].
Note that after the inverse matching, one obtains the matched GPDs $H(x,\mu_{0}=2\kappa xP_{z})$ at a varying $\overline{\rm MS}$ scale in $x$. To implement the DGLAP evolution, we set $\mu=2$ GeV and $\mu_{0}=2\kappa xP_{z}$ and obtain the evolution kernel as a triangular matrix down to a minimal value $x_{\rm min}\sim 0.15$. Then we invert the triangular matrix and apply it to the matched GPDs $H(x,\mu_{0})$, which leads to $H(x,\mu)$ at the fixed scale $\mu$ down to $x_{\rm min}\sim 0.15$.

### v2

@@ F.1 Matching coefficient
The NNLO matching kernel with LRR and RGR can be expanded to ${\cal O}(\alpha_{s})$ as Zhang:2023bxs ; Holligan:2023rex
@@
$\quad+\delta{\cal C}_{0,{\rm LRR\;rest}}{\cal C}_{\rm hybrid\;LRR}\left(\frac{x}{y},\frac{\mu_{0}}{yP_{z}},|y|\lambda_{s}\right)+{\cal O}(\alpha_{s}^{3}),$
@@
where $C_{F}=4/3$, $C_{A}=3$, $T_{F}=1/2$, and $n_{f}=3$ for our lattice ensemble.
The matching coefficient ${\cal C}$ can be derived from the NNLO kernel Li:2020xml ; Chen:2020ody in the hybrid scheme Ji:2020brr ; Gao:2021dbh ,
${\cal C}(\xi,{\mu_{0}\over p_{z}},z_{s}p_{z})$ $=\left[{\cal C}_{\rm ratio}(\xi,{\mu_{0}\over p_{z}})+{\cal C}_{\rm hybrid}(\xi,{\mu_{0}\over p_{z}},z_{s}p_{z})\right]_{+}\,,$ (29)
where $\xi=x/y$, $p_{z}=yP_{z}$, ${\cal C}_{\rm ratio}$ is the NNLO ratio scheme kernel Orginos:2017kos ; Izubuchi:2018srq , and $\left[\ldots\right]_{+}$ denotes a plus function within the domain $-\infty<\xi<\infty$. The exact expressions for ${\cal C}_{\rm ratio}$ and ${\cal C}_{\rm hybrid}$ can be found in the attached Mathematica notebook.
In the second line of eq. (23), ${\cal C}_{\rm hybrid\;LRR}$ is the LRR term in the hybrid scheme Zhang:2023bxs ; Holligan:2023rex
${\cal C}_{\rm hybrid\;LRR}(\xi,{\mu_{0}\over p_{z}},z_{s}p_{z})$ (30)
$\equiv\mu_{0}\int_{-\infty}^{\infty}{d(zp_{z})\over 2\pi}e^{-i(1-\xi)zp_{z}}(|z|-z_{s})e^{-\epsilon|z|}\theta(|z|-z_{s})$
$=-{\mu_{0}\over p_{z}}\left\{\frac{[(1-\xi)^{2}-({\epsilon\over p_{z}})^{2}]\cos[(1-\xi)z_{s}p_{z}]+2(1-x){\epsilon\over p_{z}}\sin[(1-\xi)z_{s}p_{z}]}{\pi[(1-\xi)^{2}+({\epsilon\over p_{z}})^{2}]^{2}}e^{-\epsilon z_{s}}\right\}_{+},$
The small parameter $\epsilon$ is introduced to make the Fourier transform converge, whose prescription differs from that in refs. Zhang:2023bxs ; Holligan:2023jqh by $O(\epsilon)$, which has little impact on our final results. Moreover, the matching is insensitive to the value of $\epsilon$, and in our analysis we set $\epsilon=0.002a^{-1}$. Additionally, $\delta{\cal C}_{0,{\rm LRR\;rest}}$ represents the rest LRR term
$\delta{\cal C}_{0,{\rm LRR\;rest}}$ $={\cal C}_{0,{\rm LRR}}-\alpha_{s}{\cal C}^{(1)}_{0,{\rm LRR}}-\alpha_{s}^{2}{\cal C}^{(2)}_{0,{\rm LRR}}\,,$ (31)
where
${\cal C}_{0,{\rm LRR}}$ $=N_{m}\Bigg{\{}(c_{1}+c_{2})\alpha_{s}+{c_{2}(\beta_{1}-2\beta_{0}^{2})\over 4\pi\beta_{0}}\alpha_{s}^{2}$ (32)
$\quad-\frac{e^{-\frac{2\pi}{\alpha_{s}\beta_{0}}}}{2\beta_{0}^{2}}\left[4\pi\beta_{0}+\alpha_{s}c_{1}\beta_{1}+\frac{\alpha_{s}^{2}\beta_{1}c_{2}(\beta_{1}-2\beta_{0}^{2})}{4\pi\beta_{0}}\right]\Re\left[E_{1+\frac{\beta_{1}}{2\beta_{0}^{2}}}\left(-\frac{2\pi}{\alpha_{s}\beta_{0}}\right)\right]\Bigg{\}},$
${\cal C}^{(1)}_{0,{\rm LRR}}$ $=N_{m}(1+c_{1}+c_{2}),$
${\cal C}^{(2)}_{0,{\rm LRR}}$ $=N_{m}{2\beta_{0}^{2}(1-c_{2})+\beta_{1}(1+c_{1}+c_{2})\over 4\pi\beta_{0}},$
with $E_{n}(z)=\int_{1}^{\infty}e^{-zt}/t^{n}{\rm d}t$ being the generalized exponential integral, $\Re$ denoting the real part, and
@@
with $b=\beta_{1}/(2\beta_{0}^{2})$. Note that ${\cal C}^{(1)}_{0,{\rm LRR}}$ and ${\cal C}^{(2)}_{0,{\rm LRR}}$ are the lowest two terms in the $\alpha_{s}$ expansion of ${\cal C}_{0,{\rm LRR}}$, so $\delta{\cal C}_{0,{\rm LRR\;rest}}$ is ${\cal O}(\alpha_{s}^{3})$ in perturbation theory.
Without LRR, the original hybrid matching kernel does not converge well Zhang:2023bxs . With LRR, the convergence of the matching kernel gets significantly improved, thanks to the subtraction of ${\cal C}^{(1)}_{0,{\rm LRR}}$ and ${\cal C}^{(2)}_{0,{\rm LRR}}$ terms,
${\cal C}$ $=\delta\left(\frac{x}{y}-1\right)+\frac{\alpha_{s}}{2\pi}\left[{\cal C}^{(1)}-(2\pi){\cal C}^{(1)}_{0,{\rm LRR}}{\cal C}_{\rm hybrid\;LRR}\right]$
$\qquad+\left(\frac{\alpha_{s}}{2\pi}\right)^{2}\left[{\cal C}^{(2)}-(2\pi)^{2}{\cal C}^{(2)}_{0,{\rm LRR}}{\cal C}_{\rm hybrid\;LRR}\right]+{\cal C}_{0,{\rm LRR}}{\cal C}_{\rm hybrid\;LRR}.$ (34)
In numerical implementation, we can discretize the matching kernel including the integration measure as a matrix,
${dy\over|y|}{\cal C}\left({x\over y}\right)$ $={\cal I}_{x,y}+\frac{\alpha_{s}}{2\pi}\bar{\cal C}^{(1)}_{x,y}+\left(\frac{\alpha_{s}}{2\pi}\right)^{2}\bar{\cal C}^{(2)}_{x,y}+{\cal C}^{{\rm LRR}}_{x,y}\,,$ (35)
which corresponds to eq. (F.1) term by term.
To obtain the GPD from quasi-GPD, we need to first carry out the inverse matching. With a square matching matrix which is dominated by the diagonal elements, we can easily perform a matrix inversion. However, by doing so we lose the counting of the power of $\alpha_{s}$ or the perturbative accuracy, so it would be more natural to obtain the inverse kernel by expanding in $\alpha_{s}$. Nevertheless, a naive $\alpha_{s}$-expansion of the inverse of eq. (23) would not be convergent, despite that each term has a clear hierarchy in $\alpha_{s}$ order. Instead, we should use eq. (35) and treat ${\cal I}+{\cal C}^{{\rm LRR}}\equiv{\cal J}$ as the leading term, thus the inverse kernel is
${\cal C}^{-1}$ $={\cal J}^{-1}-\frac{\alpha_{s}}{2\pi}{\cal J}^{-1}\bar{\cal C}^{(1)}{\cal J}^{-1}-\left(\frac{\alpha_{s}}{2\pi}\right)^{2}\left[{\cal J}^{-1}\bar{\cal C}^{(2)}{\cal J}^{-1}-{\cal J}^{-1}\bar{\cal C}^{(1)}{\cal J}^{-1}\bar{\cal C}^{(1)}{\cal J}^{-1}\right]\,,$ (36)
which would guarantee the convergence of perturbation theory and maintain the counting rule of perturbative accuracy.
At NLO+RGR+LRR accuracy, we truncate eq. (36) at ${\cal O}(\alpha_{s})$ and use NLO DGLAP evolution and $\beta$-function. At NNLO+RGR+LRR accuracy, we truncate eq. (36) at ${\cal O}(\alpha_{s}^{2})$ and use NNLO DGLAP evolution and $\beta$-function.
F.2 Evolution coefficient
@@
where $a_{s}=\alpha_{s}(\mu_{0})/(4\pi)$, $t=\ln(\mu^{2}_{0}/\mu^{2})$, ${\cal P}^{(0)}_{qq}$ is the 1-loop splitting function, and ${\cal P}^{V(1)}_{qq}$ and ${\cal P}^{V(2)}_{qq}$ are the 2-loop and 3-loop splitting functions for the valence quark Curci:1980uw ; Moch:2004pa . They are given by
@@
where the definitions of $P_{F}(z)$, $P_{G}(z)$, $P_{n_{f}}(z)$ and $P_{A}(z)$ can be found in ref. Curci:1980uw . As for ${\cal P}^{V(2)}_{qq}$, we take the approximate solution ${\cal P}^{V(2)}_{qq}(z)=P_{\rm ns}^{(2)-}(z)$ in eq. (4.23) of ref. Moch:2004pa .
Note that after the inverse matching, one obtains the matched GPDs $H(x,\mu_{0}=2\kappa xP_{z})$ at a varying $\overline{\rm MS}$ scale in $x$. To implement the DGLAP evolution, we set $\mu=2$ GeV and $\mu_{0}=2\kappa xP_{z}$ and obtain the evolution kernel as a triangular matrix down to a minimal value $x_{\rm min}\sim 0.15$ for $P_{z}=1.937$ GeV. Then we invert the triangular matrix and apply it to the matched GPDs $H(x,\mu_{0})$, which leads to $H(x,\mu)$ at the fixed scale $\mu$ down to $x_{\rm min}\sim 0.15$.